OpenStudy (anonymous):

Which graph represents the function of f(x) = the quantity of 4 x squared minus 4 x minus 8, all over 2 x plus 2?

OpenStudy (freckles):

\[f(x)=\frac{4x^2-4x-8}{2x+2}\] I bet these questions are similar to earlier ones and that will probably get a line with a hole

OpenStudy (freckles):

let's see what happens try to factor the top

OpenStudy (anonymous):

4x(x-x-4)=0 4x(x-1)(x+4)=0

OpenStudy (freckles):

you can not factor x from each term because there is no x in the constant term 8 you can factor 4 from each term but 8/4=2